# LoraQuant: Mixed-Precision Quantization of LoRA to Ultra-Low Bits

### v1

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across a wide range of natural language tasks (Ouyang et al., 2022; Wang et al., 2022; Zhao et al., 2023), but fine-tuning LLMs for new applications remains computationally and memory intensive. To address this challenge, low-rank adaptation (LoRA; Hu et al., 2022) has emerged as a widely adopted method for parameter-efficient fine-tuning. LoRA introduces small, task-specific low-rank matrices, and during the adaptation, only these low-rank matrices are trained while the base model is frozen.
An increasingly important use case of LoRA is LLM customization, as LLM providers (e.g., OpenAI and Google) allow users to personalize their own LLMs (OpenAI, 2025; Google Cloud, 2025). This could result in hundreds of millions of customized LLMs addressing diverse tasks, domains, and users. This imposes significant challenges of storing and using these massive customized LLMs.
A straightforward attempt to solve these challenges is to freeze the base LLM and train a separate adapter for each customization (Sheng et al., 2024). During deployment, multiple LoRAs are often loaded simultaneously due to parallel user requests, and thus, the memory footprint of LoRAs becomes a concern, especially if the GPU memory is small. This is because, although each individual LoRA is relatively lightweight, the cumulative GPU memory consumption for loading many adapters can become significant.
To scale multi-LoRA systems, Gabrielsson et al. (2024) propose a compression technique that clusters different LoRAs and enables representation sharing within each cluster. However, a key limitation of their method is that it requires recomputing the shared parameters whenever a new adapter is added. Additionally, their evaluation is primarily conducted on relatively simple tasks where the base LLM already performs well and their approach struggles on more challenging tasks.
In this paper, we propose LoraQuant, a post-training LoRA quantization method for LLM customization. Quantization is a well-established technique for compressing neural networks, and is able to substantially reduce the parameter space without degrading performance much (Frantar et al., 2023; Lin et al., 2024). Despite this progress, the quantization of LoRA has received little attention in the literature, compared with quantizing full LLMs. Although existing quantization methods can be directly applied to LoRA weights, they overlook the unique low-rank structure of LoRA and perform poorly at ultra-low precisions (e.g., 1–2 bits).
@@
In the experiments, we evaluate our method by training adapters on three representative models LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023), and Mistral-7B (Jiang et al., 2023), across diverse tasks including mathematical reasoning, code generation, and summarization. Our results demonstrate that LoraQuant achieves competitive performance even under ultra-low bitwidth for LoRA (less than 2 bits on average).
@@ 2 Related Work
Model quantization. Quantization has become an important technique for reducing the memory footprint of LLMs, enabling efficient deployment without a significant loss in accuracy. A variety of post-training methods have been proposed to compress full-precision weights into lower-bit representations. For instance, GPTQ (Frantar et al., 2023) leverages second-order information to minimize quantization error, while AWQ (Lin et al., 2024) incorporates activation statistics to guide weight quantization. InvarExplore (Wen et al., 2025) leverages model invariances (namely, rotation, scaling, and permutation) to make weights perform better after quantization. SVDQuant (Li et al., 2025) performs quantization on a weight matrix, but adds a full-precision SVD with a few dimensions to improve performance. Our work differs from SVDQuant (although also using SVD) in that we focus on LoRA quantization and use the SVD decomposition to split the LoRA into two sub-LoRAs: one containing more important information and the other less important information.
@@
Although the above approaches primarily target moderate quantization (e.g., 3–8 bits), an even more extreme direction is binarization, where weights are restricted to two values (Rastegari et al., 2016). Pure binary quantization usually achieves very low performance, and researchers have proposed mixed-precision methods where some weights are binarized while others are kept in high precision. For example, PB-LLM (Shang et al., 2024) uses an additional bit to indicate whether a weight is binarized or not, which unfortunately offsets the memory saved. BiLLM (Huang et al., 2024) also adopts mixed-precision quantization, but restricts the binarization to certain column of the weight matrix. However, BiLLM still requires an additional indicator bit, as it employs a split binarization strategy in which the weights are divided into two groups and binarized separately. The extra bit is needed to indicate the group membership of each weight.
@@
Low-rank adapter (LoRA). LoRA (Hu et al., 2022) has become a widely adopted approach for parameter-efficient fine-tuning of LLMs. Building on this idea, several extensions aim to further reduce memory overhead or improve training effectiveness. Meng et al. (2024) initialize LoRA with the SVD of the base weight matrix rather than random values, providing a stronger starting point for optimization. Zhang et al. (2023) dynamically adjust the rank of LoRA during training, allocating higher ranks to more important layers. Hao et al. (2024) and Lialin et al. (2024) address the limitation that LoRA updates may remain low-rank by iteratively merging and resampling adapters during training. Kopiczko et al. (2024) propose sharing LoRA weights across layers to reduce memory.
@@
LoRA has also been applied to fine-tuning quantized models. QLoRA (Dettmers et al., 2023) freezes the quantized base model while training an add-on LoRA in full precision. LoftQ (Li et al., 2024) and ApiQ (Liao et al., 2024) improve QLoRA adapter initialization by choosing parameters that reduce quantization errors instead of random values. QA-LoRA (Xu et al., 2024) extends QLoRA by ensuring that the adapter weights remain easily quantizable even after being merged with the original weights at the cost of reducing the representational capacity of the adapters during training.
@@
In the context of serving multiple LoRAs simultaneously, prior work has focused on both hardware improvement and LoRA compression. S-LoRA (Sheng et al., 2024) proposes a batched inference strategy to improve throughput when handling concurrent LoRA requests. Punica (Chen et al., 2024) implements a customized GPU kernel, Segmented Gather Matrix-Vector Multiplication (SGMV), which enables efficient batching across heterogeneous LoRAs. Complementary to these approaches, Gabrielsson et al. (2024) reduce the memory footprint by weight sharing among a cluster of LoRAs and assigning a few additional parameters for each LoRA. However, our experiments will show that this approach does not perform well in sophisticated tasks such as math reasoning and coding (§4.2).
@@ 3.1 Splitting a LoRA into Sub-LoRAs by SVD
For an update of neural network weights $\mathbf{W}\leftarrow\mathbf{W}+\Delta\mathbf{W}$, a low-rank adapter (LoRA; Hu et al., 2022) learns two low-rank matrices $\mathbf{B}\in\mathbb{R}^{m\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times n}$ to approximate $\Delta\mathbf{W}$, where $r\ll\min(m,n)$. In other words, the update becomes $\mathbf{W}\leftarrow\mathbf{W}+\mathbf{B}\mathbf{A}$.
@@ 3.2 Quantization Methods
After splitting a LoRA into two sub-LoRAs, we perform mixed-precision quantization. For the more important sub-LoRA $\mathbf{B}_{h}\mathbf{A}_{h}$, we use round-to-nearest quantization (RTN; Jacob et al., 2018) with higher precisions (e.g., 2 bits for a weight). For the less important sub-LoRA $\mathbf{B}_{l}\mathbf{A}_{l}$, we use binary quantization with only one bit for extreme compression. Profoundly, our mixed-precision quantization with 2 bits and 1 bit allows us to achieve less than 2 bits on average, which is a setting where existing methods cannot perform well (Shang et al., 2024; Huang et al., 2024).
@@
In implementation, we apply group-wise quantization (Jacob et al., 2018), i.e., the above procedure is performed on a group of contiguous weights instead of the entire matrix. At the cost of introducing more scaling and offset values for fine-grained treatment, this reduces quantization error.
@@
Instead, we adopt the binary quantization method in Rastegari et al. (2016), which maps values to $\{-S,+S\}$ so as to preserve more representational capacity. In other words, the quantization and dequantization processes are simply given by
@@
where $\operatorname{sign}(x)=1$ if $x\geq 0$, or 1-1 otherwise. The scaling factor $S$ is set as $S=\frac{1}{n}\|\mathbf{A}_{l}\|_{1}$, which is shown to minimize the Frobenius norm between the original weights $\mathbf{A}_{l}$ and the reconstructed one $D_{\text{bin}}(\bar{\mathbf{A}}_{l})$ (Rastegari et al., 2016). Like RTN, we also use group-wise quantization, where the scaling factor is computed separately within each group of weights.
@@ 3.3 Optimizing the Sub-LoRAs by Straight-Through Gradient Descent
In other words, we reparameterize each SVD dimension such that its quantization has a smaller error. The vector $\bm{b}_{i}^{*}$ and $\bm{a}_{i}^{*}$ are initialized with $\bm{b}_{i}$ and $\bm{a}_{i}$, respectively, and are learned through gradient-based optimization. Due to the non-differentiable dequantization function, we employ the Straight-Through Estimator (STE; Bengio et al., 2013) during backpropagation, which approximates the gradient by treating the rounding function as an identity function. This allows the gradient to flow despite the non-differentiable nature of quantization.
@@ 4.1 Settings
To evaluate the effectiveness of LoraQuant, we train LoRA for three widely used open-weight language models: LLaMA 2-7B, LLaMA 2-13B (Touvron et al., 2023), and Mistral 7B (Jiang et al., 2023). We apply the LoRAs to three distinct tasks: mathematical reasoning, code generation, and summarization. For each task, we use standard benchmark datasets for evaluation. For the mathematical reasoning task, we assess performance on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) datasets using the LM Evaluation Harness framework (Gao et al., 2024); we report pass@1 accuracy as the evaluation metric. For code generation, we evaluate our model on the HumanEval dataset (Chen et al., 2021) with the Code Generation LM Evaluation Harness framework (Ben Allal et al., 2022); the metric is the accuracy of generated programs, where a program is considered accurate if it passes all test cases. For summarization, we evaluate on the XSum dataset (Narayan et al., 2018) and report ROUGE-L (Lin, 2004) as the metric.
For LoRA adapters, we train them separately for each task, mimicking the scenario of LLM customization. For mathematical reasoning, we use the MetaMathQA dataset (Yu et al., 2023) for training, noticing that the training set of GSM8K and MATH is too small. For code generation, we train the LoRA on the Magicoder-Eval-100-Instruct dataset (Wei et al., 2024), since HumanEval is a test-only dataset. Such training and evaluation follow the common practice in prior work (Biderman et al., 2024; Meng et al., 2024). For summarization, we use the standard training split of the XSum dataset (Narayan et al., 2018).
In all experiments, we adopt a widely used LoRA setup (Biderman et al., 2024), where the rank is set to 16 and LoRA modules are inserted into every linear layer of the transformer. The training of LoRA follows that in the QLoRA study (Dettmers et al., 2023), where the base language model is quantized and frozen whereas the LoRA is trained in half precision (FP16). Training hyperparameters are provided in Appendix A. After training, we apply our LoraQuant method and other quantization baselines to the LoRA weights and evaluate the performance of each approach.
@@ 4.2 Main Results
We present the main results in Tab. 1, where we consider the following popular baselines for quantizing LoRA. GPTQ (Frantar et al., 2023) quantizes weights sequentially and adjusts the remaining weights to reduce quantization error. RTN and BIN are the standard methods introduced in §3.2. PB-LLM (Shang et al., 2024) performs mixed-precision quantization with some weight being binarized and others maintained in full precision; note that an additional indicator bit is needed for each weight. BiLLM (Huang et al., 2024) works similarly, except that each column must be either quantized or maintained in full precision while using a split binarization strategy. All baselines perform group quantization with a group size of 128, which is a common practice in the literature (Lin et al., 2024; Frantar et al., 2023).
We also consider JD-Diagonal (Gabrielsson et al., 2024), which is not a quantization method but is related to the general objective of our research: reducing memory for multiple LoRAs. Specifically, JD-Diagonal clusters different LoRAs and performs weight sharing, while introducing a few ($r$-many) additional parameters for each task. In our evaluation, we treat the three tasks as a cluster for weight sharing.
@@
We also observe that JD-Diagonal fails to achieve reasonable performance in our experiments. We hypothesize that the discrepancy between our experiment and that in Gabrielsson et al. (2024) is due to the tasks and evaluation metric. Gabrielsson et al. (2024) only consider simple tasks with in-context learning samples, which can already be handled well by the base model. In our pilot study, we find that even reducing the LoRA rank to $1$ via SVD has little impact on the performance of their trained adapters, suggesting that the LoRA parameters contribute minimally in their settings. Also, they only use ROUGE-L as their metric, while in our setting math and coding require exact matching.
@@ Appendix A Implementation Details
We follow the hyperparameter settings of Biderman et al. (2024), except that we reduce the batch size because they train across multiple GPUs while we only use a single GPU.
@@ Appendix E Limitations and Future Work
Our paper proposes a novel LoRA quantization method for LLM customization. In the experiments, we have tried four datasets and three language models (12 evaluations in total). Due to the limit of our computing resources, we are unable to perform commercial scale experiments (e.g., having millions of LoRAs). Luckily, our method treats different LoRAs independently, and thus is easily scalable, as opposed to Gabrielsson et al. (2024) who require recomputing cluster parameters. Appendix D has shown the practival values of our model, and we leave the commercial applications of our method to the industry.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across a wide range of natural language tasks (instructgpt; supernaturalinstruction; llmsurvey), but fine-tuning LLMs for new applications remains computationally and memory intensive. To address this challenge, low-rank adaptation (LoRA; lora) has emerged as a widely adopted method for parameter-efficient fine-tuning. LoRA introduces small, task-specific low-rank matrices, and during the adaptation, only these low-rank matrices are trained while the base model is frozen.
An increasingly important use case of LoRA is LLM customization, as LLM providers (e.g., OpenAI and Google) allow users to personalize their own LLMs (openai_finetune; gemini_finetune). This could result in hundreds of millions of customized LLMs addressing diverse tasks, domains, and users. This imposes significant challenges of storing and using these massive customized LLMs.
A straightforward attempt to solve these challenges is to freeze the base LLM and train a separate adapter for each customization (slora). During deployment, multiple LoRAs are often loaded simultaneously due to parallel user requests, and thus, the memory footprint of LoRAs becomes a concern, especially if the GPU memory is small. This is because, although each individual LoRA is relatively lightweight, the cumulative GPU memory consumption for loading many adapters can become significant.
To scale multi-LoRA systems, compresslora propose a compression technique that clusters different LoRAs and enables representation sharing within each cluster. However, a key limitation of their method is that it requires recomputing the shared parameters whenever a new adapter is added. Additionally, their evaluation is primarily conducted on relatively simple tasks where the base LLM already performs well and their approach struggles on more challenging tasks.
In this paper, we propose LoraQuant, a post-training LoRA quantization method for LLM customization. Quantization is a well-established technique for compressing neural networks, and is able to substantially reduce the parameter space without degrading performance much (gptq; awq). Despite this progress, the quantization of LoRA has received little attention in the literature, compared with quantizing full LLMs. Although existing quantization methods can be directly applied to LoRA weights, they overlook the unique low-rank structure of LoRA and perform poorly at ultra-low precisions (e.g., 1–2 bits).
@@
In the experiments, we evaluate our method by training adapters on three representative models LLaMA2-7B, LLaMA2-13B (llama2-7b), and Mistral-7B (mistral7b), across diverse tasks including mathematical reasoning, code generation, and summarization. Our results demonstrate that LoraQuant achieves competitive performance even under ultra-low bitwidth for LoRA (less than 2 bits on average).
@@ 2 Related Work
Model quantization. Quantization has become an important technique for reducing the memory footprint of LLMs, enabling efficient deployment without a significant loss in accuracy. A variety of post-training methods have been proposed to compress full-precision weights into lower-bit representations. For instance, GPTQ (gptq) leverages second-order information to minimize quantization error, while AWQ (awq) incorporates activation statistics to guide weight quantization. InvarExplore (invarexplore) leverages model invariances (namely, rotation, scaling, and permutation) to make weights perform better after quantization. SVDQuant (svdquant) performs quantization on a weight matrix, but adds a full-precision SVD with a few dimensions to improve performance. Our work differs from SVDQuant (although also using SVD) in that we focus on LoRA quantization and use the SVD decomposition to split the LoRA into two sub-LoRAs: one containing more important information and the other less important information.
@@
Although the above approaches primarily target moderate quantization (e.g., 3–8 bits), an even more extreme direction is binarization, where weights are restricted to two values (binnet). Pure binary quantization usually achieves very low performance, and researchers have proposed mixed-precision methods where some weights are binarized while others are kept in high precision. For example, PB-LLM (pbllm) uses an additional bit to indicate whether a weight is binarized or not, which unfortunately offsets the memory saved. BiLLM (billm) also adopts mixed-precision quantization, but restricts the binarization to certain column of the weight matrix. However, BiLLM still requires an additional indicator bit, as it employs a split binarization strategy in which the weights are divided into two groups and binarized separately. The extra bit is needed to indicate the group membership of each weight.
@@
Low-rank adapter (LoRA). LoRA (lora) has become a widely adopted approach for parameter-efficient fine-tuning of LLMs. Building on this idea, several extensions aim to further reduce memory overhead or improve training effectiveness. pissa initialize LoRA with the SVD of the base weight matrix rather than random values, providing a stronger starting point for optimization. adalora dynamically adjust the rank of LoRA during training, allocating higher ranks to more important layers. flora and relora address the limitation that LoRA updates may remain low-rank by iteratively merging and resampling adapters during training. vera propose sharing LoRA weights across layers to reduce memory.
@@
LoRA has also been applied to fine-tuning quantized models. QLoRA (qlora) freezes the quantized base model while training an add-on LoRA in full precision. LoftQ (loftq) and ApiQ (apiq) improve QLoRA adapter initialization by choosing parameters that reduce quantization errors instead of random values. QA-LoRA (qalora) extends QLoRA by ensuring that the adapter weights remain easily quantizable even after being merged with the original weights at the cost of reducing the representational capacity of the adapters during training.
@@
In the context of serving multiple LoRAs simultaneously, prior work has focused on both hardware improvement and LoRA compression. S-LoRA (slora) proposes a batched inference strategy to improve throughput when handling concurrent LoRA requests. Punica (punica) implements a customized GPU kernel, Segmented Gather Matrix-Vector Multiplication (SGMV), which enables efficient batching across heterogeneous LoRAs. Complementary to these approaches, compresslora reduce the memory footprint by weight sharing among a cluster of LoRAs and assigning a few additional parameters for each LoRA. However, our experiments will show that this approach does not perform well in sophisticated tasks such as math reasoning and coding (§4.2).
@@ 3.1 Splitting a LoRA into Sub-LoRAs by SVD
For an update of neural network weights $\mathbf{W}\leftarrow\mathbf{W}+\Delta\mathbf{W}$, a low-rank adapter (LoRA; lora) learns two low-rank matrices $\mathbf{B}\in\mathbb{R}^{m\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times n}$ to approximate $\Delta\mathbf{W}$, where $r\ll\min(m,n)$. In other words, the update becomes $\mathbf{W}\leftarrow\mathbf{W}+\mathbf{B}\mathbf{A}$.
@@ 3.2 Quantization Methods
After splitting a LoRA into two sub-LoRAs, we perform mixed-precision quantization. For the more important sub-LoRA $\mathbf{B}_{h}\mathbf{A}_{h}$, we use round-to-nearest quantization (RTN; rtn) with higher precisions (e.g., 2 bits for a weight). For the less important sub-LoRA $\mathbf{B}_{l}\mathbf{A}_{l}$, we use binary quantization with only one bit for extreme compression. Profoundly, our mixed-precision quantization with 2 bits and 1 bit allows us to achieve less than 2 bits on average, which is a setting where existing methods cannot perform well (pbllm; billm).
@@
In implementation, we apply group-wise quantization (rtn), i.e., the above procedure is performed on a group of contiguous weights instead of the entire matrix. At the cost of introducing more scaling and offset values for fine-grained treatment, this reduces quantization error.
@@
Instead, we adopt the binary quantization method in binnet, which maps values to $\{-S,+S\}$ so as to preserve more representational capacity. In other words, the quantization and dequantization processes are simply given by
@@
where $\operatorname{sign}(x)=1$ if $x\geq 0$, or 1-1 otherwise. The scaling factor $S$ is set as $S=\frac{1}{n}\|\mathbf{A}_{l}\|_{1}$, which is shown to minimize the Frobenius norm between the original weights $\mathbf{A}_{l}$ and the reconstructed one $D_{\text{bin}}(\bar{\mathbf{A}}_{l})$ (binnet). Like RTN, we also use group-wise quantization, where the scaling factor is computed separately within each group of weights.
@@ 3.3 Optimizing the Sub-LoRAs by Straight-Through Gradient Descent
In other words, we reparameterize each SVD dimension such that its quantization has a smaller error. The vector $\bm{b}_{i}^{*}$ and $\bm{a}_{i}^{*}$ are initialized with $\bm{b}_{i}$ and $\bm{a}_{i}$, respectively, and are learned through gradient-based optimization. Due to the non-differentiable dequantization function, we employ the Straight-Through Estimator (STE; STE) during backpropagation, which approximates the gradient by treating the rounding function as an identity function. This allows the gradient to flow despite the non-differentiable nature of quantization.
@@ 4.1 Settings
To evaluate the effectiveness of LoraQuant, we train LoRA for three widely used open-weight language models: LLaMA 2-7B, LLaMA 2-13B (llama2-7b), and Mistral 7B (mistral7b). We apply the LoRAs to three distinct tasks: mathematical reasoning, code generation, and summarization. For each task, we use standard benchmark datasets for evaluation. For the mathematical reasoning task, we assess performance on the GSM8K (gsm8k) and MATH (minervamath) datasets using the LM Evaluation Harness framework (lmeval); we report pass@1 accuracy as the evaluation metric. For code generation, we evaluate our model on the HumanEval dataset (humaneval) with the Code Generation LM Evaluation Harness framework (bigcode); the metric is the accuracy of generated programs, where a program is considered accurate if it passes all test cases. For summarization, we evaluate on the XSum dataset (xsum) and report ROUGE-L (rouge) as the metric.
For LoRA adapters, we train them separately for each task, mimicking the scenario of LLM customization. For mathematical reasoning, we use the MetaMathQA dataset (metamathqa) for training, noticing that the training set of GSM8K and MATH is too small. For code generation, we train the LoRA on the Magicoder-Eval-100-Instruct dataset (magicoder), since HumanEval is a test-only dataset. Such training and evaluation follow the common practice in prior work (learnless; pissa). For summarization, we use the standard training split of the XSum dataset (xsum).
In all experiments, we adopt a widely used LoRA setup (learnless), where the rank is set to 16 and LoRA modules are inserted into every linear layer of the transformer. The training of LoRA follows that in the QLoRA study (qlora), where the base language model is quantized and frozen whereas the LoRA is trained in half precision (FP16). Training hyperparameters are provided in Appendix A. After training, we apply our LoraQuant method and other quantization baselines to the LoRA weights and evaluate the performance of each approach.
@@ 4.2 Main Results
We present the main results in Tab. 1, where we consider the following popular baselines for quantizing LoRA. GPTQ (gptq) quantizes weights sequentially and adjusts the remaining weights to reduce quantization error. RTN and BIN are the standard methods introduced in §3.2. PB-LLM (pbllm) performs mixed-precision quantization with some weight being binarized and others maintained in full precision; note that an additional indicator bit is needed for each weight. BiLLM (billm) works similarly, except that each column must be either quantized or maintained in full precision while using a split binarization strategy. All baselines perform group quantization with a group size of 128, which is a common practice in the literature (awq; gptq).
We also consider JD-Diagonal (compresslora), which is not a quantization method but is related to the general objective of our research: reducing memory for multiple LoRAs. Specifically, JD-Diagonal clusters different LoRAs and performs weight sharing, while introducing a few ($r$-many) additional parameters for each task. In our evaluation, we treat the three tasks as a cluster for weight sharing.
@@
We also observe that JD-Diagonal fails to achieve reasonable performance in our experiments. We hypothesize that the discrepancy between our experiment and that in compresslora is due to the tasks and evaluation metric. compresslora only consider simple tasks with in-context learning samples, which can already be handled well by the base model. In our pilot study, we find that even reducing the LoRA rank to $1$ via SVD has little impact on the performance of their trained adapters, suggesting that the LoRA parameters contribute minimally in their settings. Also, they only use ROUGE-L as their metric, while in our setting math and coding require exact matching.
@@ Appendix A Implementation Details
We follow the hyperparameter settings of learnless, except that we reduce the batch size because they train across multiple GPUs while we only use a single GPU.
@@ Appendix E Limitations and Future Work
Our paper proposes a novel LoRA quantization method for LLM customization. In the experiments, we have tried four datasets and three language models (12 evaluations in total). Due to the limit of our computing resources, we are unable to perform commercial scale experiments (e.g., having millions of LoRAs). Luckily, our method treats different LoRAs independently, and thus is easily scalable, as opposed to compresslora who require recomputing cluster parameters. Appendix D has shown the practival values of our model, and we leave the commercial applications of our method to the industry.